Intro to physics help please!

A sleeping 68 kg man has a metabolic power of 75 W
Answer #1

Solution:

Energy used over the time of t = 8 hours

t = (8 h)*(60 min / 1 h)*(60 sec / 1 min)

t = 28800 s (8 hours has 28800 seconds)

The metabolic power of the man is,

P = 75 W

P = 75 J/s (1 watt = 1 joule/sec)

The relationship between the power P, energy E and time t is given by,

P = E/t

E = P*t

E = (75 J/s)*(28800 s)

E = 2.16*106 J

We have, 1 Cal = 4184 Joules

E = (2.16*106 J)*(1 Cal / 4184 J)

E = 516.25 Cal

The person burns 516.3 Cal in 8 hours of sleep.

Hence the answer is 516.3 Cal